# Elucidating the representation of images within an unconditional diffusion model denoiser

### v1

@@ 3 Stochastic image reconstruction by conditional sampling
In this section, we propose an algorithm to “decode” $\phi$. As with any other representation, decoding $\phi$ is essential to understanding what it represents. Computing $\phi$ involves taking spatial average of the activations, that renders the representation translation invariance. Hence, $\phi$ represents not only the image it is computed from, but a whole set of images which have different activations with the same spatial average. In other words, $\phi$ is a many to one function. We aim to sample from the density embedded within the denoiser (or equivalently the score network) using a diffusion algorithm, while requiring that all samples share the same $\phi(x^{c})$ computed from a target image, $x^{c}$. This is a conditional sampling solutions for a non-linear inverse problem. The forward model, however, is not an auxiliary network or external measurement function, but the down-sampling (Encoding) path of the same score network. Since the algorithm samples images that are consistent with the network’s own representation of a target conditioner image, we view it as implementing a stochastic reconstruction from the representation.
The algorithm is obtained by augmenting a reverse diffusion algorithm with a soft projection onto the non-convex set in pixel space defined by $\phi$. To achieve this, the score step is alternated with an iterative projection step, where the sample is modified until its representation matches $\phi(x^{c})$, analogous to methods used to generate texture images from their measured statistics [17]. The matching is implemented by minimizing the Euclidean distance between the sample and target representations, $\|\phi(x)-\phi(x^{c})\|^{2}$, which is achieved by back-propagating the gradient of the loss through the first half of the network (i.e. $\bar{a}_{4}(x_{\sigma})$). The algorithm, hence, consists of two alternating steps described in Algorithms 2 and 1: At every time point in the synthesis, first the $\phi$ of the sample is matched to the target’s, via back propagating a gradient in the score network. Second the score-directed step pushes the sample closer to $p(x)$ by removing noise.
@@
Draw $x_{T}\sim\mathcal{N}(m,\sigma_{T}^{2}\mathrm{Id})$triangleright Initialize sample
@@
    $x_{t}\leftarrow$ Guidance step $(x_{t},x_{t}^{c})$triangleright Update $x_{t}$ to match representations using Algorithm 2
@@
    $x_{t}\leftarrow x_{t}+s(x_{t})+\sigma_{t-1}z$triangleright Update $x_{t}$ using the score
@@
$\phi_{x^{c}_{t}}=\phi(x^{c}_{t})\hskip 5.69046pt\text{and}\hskip 5.69046pt\phi_{x_{t}}=\phi(x_{t})$triangleright Compute representations
$\mathcal{L}(x_{t})=\|\phi(x_{t})-\phi_{x^{c}_{t}}\|^{2}$triangleright Compute distance between representations
@@
    $\nabla_{x_{t}}\mathcal{L}(x_{t})=(\phi_{x^{c}_{t}}-\phi(x_{t}))^{T}\nabla_{x_{t}}\phi(x_{t})$triangleright Compute gradient
    $x_{t}\leftarrow x_{t}-\eta\nabla_{x_{t}}\mathcal{L}(x_{t})$triangleright Update
@@
This procedure can be described as a guided sampling algorithm [23], where the synthesis is guided by the network’s own representation. Matching the $\phi$’s changes the trajectory by forcing the sample to be within the set of images whose $\phi$ is equal to the conditioner’s. For a visualization of the effect matching on $x_{\sigma}$ in one step see Figures 11 and 12. The matching step does not change the noise level (Figure 13), so it is in a sense "orthogonal" to the score step. This means that matching iterations guide the sample in the domain of images that are high probability according to $p_{\sigma}(x_{\sigma})$ to reach the set of images defined by $\phi(x_{\sigma}^{c})$.
@@
Figures 3 and 10 show samples generated by the algorithm. In each panel, the representation is obtained from an image selected from the validation set of ImageNet64 [24]. Then the stochastic reconstruction algorithm is used to generate $8$ images with the same $\phi$. The conditionally generated samples are not identical, but all are visually similar to the conditioner image $x^{c}$. Importantly, they share location-specific global structure as well as location-nonspecific details. Thus, the visual commonalities and diversities in the samples reveal what is and is not captured by the spatial averages of the feature vector in the middle block.
Figure 3: Samples from the stochastic reconstruction algorithm 1. Image at center of each panel is a target image drawn from the validation set of ImageNet64. The surrounding eight images are reconstruction samples, conditioned on the representation vector $\phi$ obtained from the target image. Samples in each panel are semantically similar. The diversity of samples varies for different target images, demonstrating that the sets defined by different $\phi$’s are drawn from neighborhoods of different sizes. More examples in Appendix B.
4 Properties of the representation
In the previous sections, we extracted a sparse representation from the output layer of the middle block and developed a stochastic algorithm to recover an image from it. In this section, we explore some properties of $\phi$. In particular, we examine the components of this vector, and their relationship to the content of the (clean) image $x$. Then, we study semantic similarities in pixel space in relation to Euclidean distances in the representation space.
4.1 Specialized versus common channels
For each channel, $i$, in the output layer of the middle block we quantify its (non-)selectivity using the participation ratio:
$\frac{\|c(i)\|_{1}^{2}}{n\ \|c(i)\|_{2}^{2}}$
where $c(i)=\left(\phi({x_{1}}_{\sigma})[i],...,\phi({x_{n}}_{\sigma})[i]\right)$ is the concatenation of the $i^{th}$ entry of $\phi$ for all images in the ImageNet64 validation set (approximately $50k$ images). This provides an estimate of the number of active channels, with smaller values indicating more selectivity. The threshold is set according to activation at zero noise level (Details in Appendix C). Figure 4 shows the distribution of selectivity for all $512$ channels in the middle block. The distribution is bimodal. Channels fall roughly into one of the two categories: selective channels that respond only to a small fraction of images in the set, and non-selective channels that respond to most images.
Figure 4: Channel selectivity. Left: Participation ratios for each channel over ImageNet. Distribution is bimodal, correponding to channels that are highly specialized (and infrequently active) on left, and commonly used on right. Right: The panels show the set of images that maximally activate each of four particular specialized channels, revealing selectivity for rectangular periodic lattices, a bird on a branch, full-frame cylindrical objects, and dog faces. Sorted by selectivity $i={87,143,187,287}$.
Specialized channels respond to specific features or patterns, and have negligible responses for most images in the dataset [25]. Figure 4 shows sets of images that maximally excite of four example specialized channels. Images within each group clearly share similar patterns. The observed sparsity of $\phi$ is a consequence of the existence of specialized channels. Since each image only contains a subset of patterns out of all the patterns present in the data distribution, only a handful of specialized channels are activated for each image. Figure 5 shows an example image along with the top three specialized channels it activates. Each channel is selective for a particular property of the image. Over the entire dataset, all channels are used, but different channels are used for different images.
The interpretation of the non-selective (“common”) channels is less straightforward. They are on average larger in magnitude, respond to a larger set of frequencies and orientations, and are the most stable w.r.t to noise level. These imply that they capture shared or common image features, such as brightness, global structure of a scene, etc.
4.2 Distances in representation space
An immediate consequence of sparsity and channel selectivity is that the $\phi$’s lie in a union of subspaces in $\mathbb{R}^{d}$, whose dimensionality lies approximately in the range $[0.2d,0.3d]$ (see Figure 1). How many of these subspaces are used? On one extreme, if the zero entries of $\phi$ coincide for all images $x$, this number could be as small as one. On the other extreme, it is bounded above by the number of possible $k$-dimensional subspaces, approximately $\sum_{k=100}^{150}\binom{500}{k}$, which far exceeds the size of any feasible dataset. The assumption behind this extreme case is that features corresponding to entries of $\phi$ are independent from one another, so that all combinations of features are equally likely. We know intuitively that this is not the case. Only certain combinations are likely and frequent in the training data distribution. This implies that collections of $\phi$’s should lie in common subspaces (Figure 6).
The network is trained solely with a denoising objective, without labels, augmentations, or regularization to induce any specific type of similarity or grouping (e.g membership in the same class). Therefore proximity in the representation space defines a fully unsupervised form of similarity in the image space. What do images whose $\phi$’s lie on the same subspace have in common? To examine this, we gathered images from the dataset with highest cosine similarity to a given target image in the representation space. Figure 7 shows two sets of those images. Images that are closest in the $\phi$ are visually and semantically similar, in stark contrast to images that are closest in the pixel space.
@@ 4.2 Distances in representation space
The observation that Euclidean distances in representation space are related to perceptual similarities in the image space can be expressed more precisely. Each $\phi$ characterizes a conditional density, ${p}(x|\phi)$ from which Algorithm 1 draws samples. Distances between these conditional densities should govern the separation of $\phi$’s. We show that Euclidean distances between $\phi$’s are correlated with distances between conditional densities they induce. This is captured by a Euclidean embedding property, which ensures that the separation of $\phi$’s is related to a distance between the probability distributions ${p}(x|\phi)$, and hence that there exists $0<A\leq B$ with $B/A$ not too large, such that
$\forall x_{1},x_{2}\leavevmode\nobreak\ \leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ A\|\phi(x_{1})-\phi(x_{2})\|^{2}\leq d^{2}({p}_{1},{p}_{2})\leq B\|\phi(x_{1})-\phi(x_{2})\|^{2}.$ (3)
@@
where ${p}_{1}={p}_{\sigma}(x_{\sigma}|\phi_{1})$ and ${p}_{2}={p}_{\sigma}(x_{\sigma}|\phi_{2})$. This distance is based on the difference in the expected score assigned to $x_{\sigma}$ by ${p}(x_{\sigma}|\phi_{1})$ versus ${p}(x_{\sigma}|\phi_{2})$, integrated across all noise levels. It provide a distance by symmetrizing the Kullback-Leibler divergence $\text{KL}(p\|q)$ between two distributions $p$ and $q$ proved in [26]:
[…]
The rightmost panel in Figure 7 confirms that, for a set of randomly selected pairs of images, Euclidean distances of their representations satisfy the Euclidean embedding inequality (3) with $\frac{B}{A}=2.9$.
4.3 Emergent clusters
Emergent clusters Human-labeled object classes Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 8: Representation vectors, $\phi$, are concentrated within emergent clusters, less so with respect to object classes. Top Left: Euclidean distance between centroids of two clusters plotted against square root of their average variance along the line connecting the centroids, for 100 randomly selected cluster pairs. For all pairs, centroids are more than two standard deviations apart, and thus well-separated. t-SNE visualization of 10 randomly selected clusters. Top Right: Analogous plots for pairs of human-labeled classes, which are substantially more overlapping. Bottom: Each row contains images randomly selected from a cluster. Images within a cluster are visually similar and share global organization and semantic patterns, but they not necessarily from the same class. For example, dogs are clustered based on their pose, but not their breed. Thus, from a learned unconditional denoiser, a new notion of similarity emerges. More examples in Appendix C.
The fact that pairwise Euclidean distances in the $\phi$ space are meaningful implies that similar images cluster together within the subspace. To test this, we applied a K-Means clustering algorithm [27] to the $\phi$ vectors computed from the ImageNet dataset (See Appendix C). Figure 8 shows that within clusters the $\phi$’s are well concentrated around centroids, and the centroids are well separated: centroids of adjacent clusters are at least 2 standard deviations apart. The $\phi$’s of human labeled classes, by comparison, are less well separated. As a result, we expect images from different classes to appear within the same cluster. Examples are shown in Figure 8 for 9 clusters (see Appendix C for more).
Images within same clusters clearly share location-specific global scene structure, as well as location-nonspecific detailed elements (e.g., objects and their constituent parts, texture). However, in many cases, the similarity is not the object identity: a variety of circular objects appear in one cluster, athletes or dogs on a grass background in another, and a variety of sea creatures on a blue background in another. What connects images within the same cluster is "the gist of the scene" [28, 29, 30, 31, 32], which is only partially consistent with human labeled class memberships in ImageNet. Remarkably, these emergent clusters arise solely from learning the score, by minimizing a denoising objective. In the unconditional model, there is no external source of information to dictate proximity in the latent space. Importantly, this reveals that the type of patterns that are useful for computing the score are not object identities.
@@ 5 Discussion
Generative diffusion models have shown incredible success in learning and sampling from image densities. This feat is due to near minimum mean square error of denoising networks used in these algorithms. We hypothesized that these networks must construct an internal representation of image features that differentiate signal from noise, and we sought to elucidate that representation. We found that the vector of spatial averages of channels in the middle block of a trained UNet can provide a signature of this representation. Across a dataset of photographic images, the components of this vector are sparse – a relatively small subset of channels are active for each image – and thus the scores of complex image distributions have a low-dimensional structure, which is made explicit in these networks.
In addition to the sparsity of the representation vectors, we found that the geometry of the representation space is meaningful. Images with similar representations are visually similar and share location-specific global structure and location-nonspecific details. We developed a stochastic reconstruction algorithm, and showed that Euclidean distances in representation space are correlated with the symmetrized KL divergence of reconstruction distributions. Finally, we found that unsupervised clustering of the representation vectors yields well-separated groups of images that are semantically related, but only partially aligned with object identity. Thus, a network trained “bottom-up”, using only a denoising objective and no external conditioning, labelling, augmentation or regularization, can nevertheless learn a rich and accessible representation of image structure.
@@ A.1 Architecture
We used the original UNet architecture [5] with a few modifications. The architecture is made up of 3 encoder blocks, 1 middle block, and 3 decoder blocks. The number of layers in each block is 2, 3, and 6 for the encoder, middle and decoder blocks. Each layer consists of $3\times 3$ convolutions with zero boundary handling, Layer normalization and ReLU. The number of channels is 64 in the first encoder block and then grows by a factor of two after each downsampling operation, and decreased by a factor of two after each upsampling operation. This results in 64, 128, 256 channels in the encoder blocks, 512 channels in the middle block, and 256, 128, 64 in the decoding blocks. The total number of parameters is $\sim 13m$.
The receptive field size at the end of the middle block is $84\times 84$ surpassing the spatial size of the input images in the ImagNet dataset. This is required for capturing large structures in the absence of attention blocks [19]. The depth of decoder blocks is increased in order to increase the expressivity of the denoising operations. This is again related to the size of the receptive field at the end of each decoder block with respect to the representation at the end of the middle block.
Figure 9: Modified UNet architecture used in our experiments.
A.2 Dataset
We used down-sampled version of ImageNet data set [24] , 3×,64×643$\times$,64$\times$ 643 × , 64 × 64, throughout the paper. The training set consists of $\sim 1.2m$ images and the validation set consists of $\sim 50k$ images, of objects, animals, scenes, etc. We did not use the class labels for training.

### v2

@@ 3 Stochastic image reconstruction by conditional sampling
In order to define the representation more precisely, we propose an algorithm to “decode” $\phi$. As with any other representation, decoding $\phi$ is essential to exposing what it represents. Computing $\phi$ involves taking spatial average of the activations. Hence, $\phi$ represents not only the image it is computed from, but a whole set of images which have different activations with the same spatial average. In other words, $\phi$ is a many to one function. The goal is to sample from $p(x)$ using a diffusion algorithm, while requiring that all samples have the same $\phi(x^{c})$ computed from a target image, $x^{c}$. Since the algorithm samples images that are consistent with the network’s own representation of a target image, it is a stochastic reconstruction from the representation.
The algorithm is obtained by augmenting a reverse diffusion algorithm with a soft projection onto the non-convex set in pixel space defined by $\phi$. To achieve this, the score step is alternated with an iterative projection step, where the sample is modified until its representation matches $\phi(x^{c})$. This matching step is analogous to methods used to generate texture images from their measured statistics (Portilla and Simoncelli, 2000). The matching is implemented by minimizing the Euclidean distance between the sample and target representations, $\|\phi(x)-\phi(x^{c})\|^{2}$, which is achieved by back-propagating the gradient of the loss through the first half of the network (i.e. $\bar{a}_{4}(x_{\sigma})$). The algorithm, hence, consists of two alternating steps described in Algorithms 2 and 1: At every time point in the synthesis, first the $\phi$ of the sample is matched to the target’s, via back propagating a gradient in the score network. Then, the score-directed step pushes the sample closer to $p(x)$ by removing noise via using the decoder of the UNet.
@@
Draw $x_{T}\sim\mathcal{N}(m,\sigma_{T}^{2}\mathrm{Id})$ $\triangleright$ Initialize sample
@@
  $x_{t}\leftarrow$ Guidance step $(x_{t},x_{t}^{c})$ $\triangleright$ Update $x_{t}$ to match representations using Algorithm 2
@@
  $x_{t}\leftarrow x_{t}+s(x_{t})+\sigma_{t-1}z$ $\triangleright$ Update $x_{t}$ using the score
@@
$\phi_{x^{c}_{t}}=\phi(x^{c}_{t})\hskip 5.69046pt\text{and}\hskip 5.69046pt\phi_{x_{t}}=\phi(x_{t})$ $\triangleright$ Compute representations
$\mathcal{L}(x_{t})=\|\phi(x_{t})-\phi_{x^{c}_{t}}\|^{2}$ $\triangleright$ Compute distance between representations
@@
  $\nabla_{x_{t}}\mathcal{L}(x_{t})=(\phi_{x^{c}_{t}}-\phi(x_{t}))^{T}\nabla_{x_{t}}\phi(x_{t})$ $\triangleright$ Compute gradient
  $x_{t}\leftarrow x_{t}-\eta\nabla_{x_{t}}\mathcal{L}(x_{t})$ $\triangleright$ Update
@@
This procedure can be described as a "self-guided" sampling algorithm , where the synthesis is guided by the network’s own representation. Matching the $\phi$’s changes the trajectory by forcing the sample to be within the set of images whose $\phi$ vectors are equal to the conditioner’s. The matching step does not change the noise level (Figure 20), so it is in a sense "orthogonal" to the score step. This means that matching iterations guide the sample in the domain of images that are high probability according to $p_{\sigma}(x_{\sigma})$ to reach the set of images defined by $\phi(x_{\sigma}^{c})$. A conceptual diagram of the algorithm is shown in Figure˜8.
@@
Figure 7 shows samples generated by the algorithm. In each panel, the representation is obtained from a real image. Then the stochastic reconstruction algorithm is used to generate $8$ images with the same $\phi$. The conditionally generated samples are not identical, but all are visually similar to the conditioner image $x^{c}$. Importantly, they share location-specific global structure as well as location-nonspecific details. Thus, the visual commonalities and diversities in the samples reveal what is and is not captured by the spatial averages of the feature vector in the middle block.
Figure 7: Samples from the stochastic reconstruction algorithm 1, using models trained unconditionally on Texture, LSUN-Bedrooms, ImageNet64, and CelebA datasets (from top to bottom row). Image at center of each panel is an original image from which the target $\phi$ is computed. The surrounding eight images are samples, conditioned on that $\phi$. Semantic similarities between samples within each panel reveal image structures captured by $\phi$. See Appendix C for more examples and also samples from the same models without conditioning.
@@
Each $\phi$ characterizes a conditional density, ${p}(x|\phi)$, from which Algorithm 1 draws samples. We show that Euclidean distances between $\phi$’s are correlated with distances between conditional densities they induce. This is captured by a Euclidean embedding property, which ensures that the separation of $\phi$’s is related to a distance between the probability distributions ${p}(x|\phi)$, and hence that there exists $0<A\leq B$ with $B/A$ not too large, such that
$\forall x_{1},x_{2}\penalty 10000\ \penalty 10000\ ,\penalty 10000\ \penalty 10000\ A\|\phi(x_{1})-\phi(x_{2})\|^{2}\leq d^{2}({p}_{1},{p}_{2})\leq B\|\phi(x_{1})-\phi(x_{2})\|^{2}.$ (3)
@@
where ${p}_{1}={p}_{\sigma}(x_{\sigma}|\phi_{1})$ and ${p}_{2}={p}_{\sigma}(x_{\sigma}|\phi_{2})$. This distance is based on the difference in the expected score assigned to $x_{\sigma}$ by ${p}(x_{\sigma}|\phi_{1})$ versus ${p}(x_{\sigma}|\phi_{2})$, integrated across all noise levels. It provide a distance by symmetrizing the Kullback-Leibler divergence $\text{KL}(p\|q)$ between two distributions $p$ and $q$ proved in (Verdú, 2010; Lyu, 2012):
$\text{KL}{(p\|q)}=\int_{0}^{\infty}\mathbb{E}_{p_{\sigma}}[\|\nabla_{x_{\sigma}}\log{p_{\sigma}}(x_{\sigma})-\nabla_{x_{\sigma}}\log q_{\sigma}(x_{\sigma})\|^{2}]\,\sigma d\sigma.$
The right panel of Figure˜8 confirms that, for a set of random pairs of images, Euclidean distances of their representations satisfy the Euclidean embedding inequality (3) with $\frac{B}{A}=2.9$. See fig.˜21 for an intuition about what the symmetrized distance measures.
Figure 8: Left: Schematic of alternating matching and score steps. Right The distance between random pairs of $\phi$ vectors is strongly related to the distance between the pair of conditional densities they induce.
@@ 4 Discussion
Generative diffusion models have shown incredible success in learning and sampling from image densities. This feat is due to near minimum mean square error of denoising networks used in these algorithms. We hypothesized that these networks must construct an internal representation of image features that differentiate signal from noise, and we sought to elucidate that representation. We found that the vector of spatial averages of channels in the middle block of a trained UNet can provide a signature of this representation. The components of this vector are sparse and thus the scores of complex image distributions have a low-dimensional structure, which is made explicit in these networks. We also found that the geometry of the representation space is meaningful and nearby images in representation space are semantically similar in pixel space. We showed that unsupervised clustering of the representation vectors yields well-separated groups of images that are semantically related, but only partially aligned with object identity. We developed a stochastic reconstruction algorithm, and showed that Euclidean distances in representation space are correlated with the symmetrized KL divergence of reconstruction distributions. These results show that a network trained “bottom-up”, using only a denoising objective and no external conditioning, labeling, augmentation or regularization, can nevertheless learn a rich and accessible representation of image structure.
Limitations and future directions. Many open questions remain to be explored. One concerns a deeper understanding of the geometry of the representation space. We observed that the denoiser transforms a union of manifolds into a union of subspaces. What is the distribution of the dimensionalities of these subspaces, and how many of them, out of all possible ones, are spanned by the $\phi$’s? Importantly, can we estimate the joint distribution of $\phi$’s within these subspaces? A more comprehensive understanding of the latent space could be used to generalize Algorithm 1 to other conditional settings—for example, sampling from an emergent cluster given its centroid, or combining features to create new subspaces, thereby enabling out-of-distribution generalization. Another direction is to investigate the role of the encoder and decoder blocks in representing image structure, and how these representations depend on the noise level. This requires analyzing the interaction between layer depth and noise level, a relationship that is complex but essential to characterize. Finally, while our analysis has focused on fully convolutional neural networks, it is important to examine whether similar representations with comparable properties also arise in more modern architectures such as transformers.
@@ A.1 Architecture
We used the original UNet architecture [Ronneberger et al., 2015] with a few modifications. The architecture is made up of 3 encoder blocks, 1 middle block, and 3 decoder blocks. The number of layers in each block is 2, 3, and 6 for the encoder, middle and decoder blocks. Each layer consists of $3\times 3$ convolutions with zero boundary handling, Layer normalization and ReLU. The number of channels is 64 in the first encoder block and then grows by a factor of two after each downsampling operation, and decreased by a factor of two after each upsampling operation. This results in 64, 128, 256 channels in the encoder blocks, 512 channels in the middle block, and 256, 128, 64 in the decoding blocks. The total number of parameters is $\sim 13m$.
The receptive field size at the end of the middle block is $84\times 84$ surpassing the spatial size of the input images in the ImagNet dataset. This is required for capturing large structures in the absence of attention blocks [Kadkhodaie et al., 2023]. The depth of decoder blocks is increased in order to increase the expressivity of the denoising operations. This is again related to the size of the receptive field at the end of each decoder block with respect to the representation at the end of the middle block.
Figure 9: Fully convolutional UNet architecture used in our experiments.
A.2 Datasets
We show results from models trained on 4 public datasets:
  • ImageNet64: down-sampled version of ImageNet data set [Deng et al., 2009] , $3\times 64\times 64$. The training set consists of $\sim 1.2m$ images and the validation set consists of $50k$ images, of objects, animals, scenes, etc. We did not use the class labels for training.
  • LSUN-Bedroom dataset [Yu et al., 2015]: down-sampled to $3\times 80\times 80$ images. We trained the model on a subset of images randomly selected. Training set size $\sim 300,000$ images.
  • CelebA dataset [Liu et al., 2015]: down-sampled to $3\times 80\times 80$ images. Training set size $\sim 200,000$ images.
  • Texture dataset (collected by the authors, not published), cropped to $3\times 80\times 80$ images. Training set size $\sim 230,000$ images.